Let {an} be a sequence that satisfies the recurrence relation an=anà ƒ ƒ ¢ ˆ ’1+3anà ƒ ƒ ¢ ˆ ’2, for n=2,3,.,.,., where a0=1,a1=2. Find the values of a2,a3.
1
Expert's answer
2021-10-13T17:38:06-0400
"a_n=a_{n-1}+3a_{n-2}, n\geq2"

"a_0=1, a_1=2"


"a_2=a_1+3a_0=2+3(1)=5"

"a_3=a_2+3a_1=5+3(2)=11"